OpenStudy (anonymous):

Let f(7)=0, f'(7)=14, g(7)=1, g'(7)=1/7. Find h'(7) if h(x)=f(x)/g(x).

zepdrix (zepdrix):

So they want you to get an understanding of the product rule with this one.\[\huge h'(x)=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^2}\]Understand what's going on? :) From here, you just fill in the pieces that they gave you.

OpenStudy (anonymous):

So the answer is 14?

zepdrix (zepdrix):

\[\huge h'(x)=\frac{14\cdot 1-0\cdot\frac{1}{7}}{\left(\frac{1}{7}\right)^2}\] I think you get this when you plug everything in. Which ends up being.... 14*49 if I did the calculations correctly.

OpenStudy (anonymous):

I thought that the denominator is 1, not 1/7.

zepdrix (zepdrix):

ah good call :) lol So 14, yes. good job \^^/

OpenStudy (anonymous):

Thank you!